Consider the following economic data:

Unemployment rate= 5%
Inflation rate= 2%
Real GDP= $18 trillion
Growth rate= 2%
Natural rate of unemployment= 4%
Potential GDP= $20 trillion
Marginal Propensity to consume=.8

Based on the data, answer the following:

1. What data indicate there is a problem? Explain.

2. Identify what kind of GDP gap this economy is experiencing. Why is this a problem?

3. State a fiscal policy option to fix the problem. How will your proposal fix the problem?

4. State a monetary policy option to fix the problem. How will your proposal fix the problem?

5. Given the power of the multiplier, how much must spending change to reach potential GDP?
1
Expert's answer
2016-04-26T08:38:06-0400
Unemployment rate = 5%
Inflation rate = 2%
Real GDP = $18 trillion
Growth rate = 2%
Natural rate of unemployment = 4%
Potential GDP = $20 trillion
Marginal Propensity to consume = 0.8

1. There is a problem, because potential GDP is higher than real and there is 1% cyclical unemployment.
2. There is a recessional GDP gap in this economy, because real GDP is below potential GDP.
3. The expansionary fiscal policy option (decrease in taxes) should be implemented to fix the problem. In this case consumption will increase and real GDP will increase too to the level of potential GDP.
4. An expansionary monetary policy option (increase in money supply) should be implemented to fix the problem. In this case consumption will increase and real GDP will increase too to the level of potential GDP.
5. Multiplier equals m = 1/(1 - c) = 1/(1 - 0.8) = 5
Given the power of the multiplier, spending should be increased by (20 - 18)/5 = $0.4 trillion to reach potential GDP.